OpenStudy (princesssleelee):

Use the graph below to answer the following question: graph of parabola going through negative 1, 1, and 1, negative 3 What is the average rate of change from x = –1 to x = 1? –4 –2 2 4

OpenStudy (princesssleelee):

@mathmale @mathmate

OpenStudy (princesssleelee):

@igigighjkl

OpenStudy (igigighjkl):

Sup.

OpenStudy (igigighjkl):

Do you want awnser

OpenStudy (princesssleelee):

yessss

OpenStudy (igigighjkl):

This might be correct your coordinates are 0,3 and 2,-5 -5-3 = -8 2-0 = 2 -8 divided by 2= -4

OpenStudy (igigighjkl):

Was this helpful?

OpenStudy (princesssleelee):

possibly, do you know of anyone who you could tag to double check it?

OpenStudy (igigighjkl):

None of my fans (exept you) is here. sorry.

OpenStudy (igigighjkl):

Are you on Flvs?

OpenStudy (princesssleelee):

its okay ill take your word :) ill let you know if youre correct. are you familiar with this?6

OpenStudy (igigighjkl):

Summon me after.

OpenStudy (princesssleelee):

ok

OpenStudy (igigighjkl):

hello?

OpenStudy (princesssleelee):

im here

OpenStudy (princesssleelee):

Of the following, which is the solution to 2x2 + 3x = –5? the quantity negative three plus or minus i times the square root of thirty one over four the quantity negative three plus or minus i times the square root of thirty one over two the quantity three plus or minus i times the square root of thirty one over four the quantity negative three plus or minus i times the square root of fifty over four

OpenStudy (princesssleelee):

@AlexandervonHumboldt2

OpenStudy (princesssleelee):

@mathstudent55

OpenStudy (igigighjkl):

Sorry I was doing something I have to go.

OpenStudy (princesssleelee):

ok